# Muon Optimizes Under Spectral Norm Constraints

### v1

@@ 6.2 Generalizations of Muon
Quadratic form $\textup{Tr}(\mathbf{X}^{\top}\mathbf{M}\mathbf{X})$ All matrices ($\mathbf{M}\succeq\mathbf{0}$) Smooth quadratic form, convex
@@
More generally, the constrained optimization problem $\min_{\mathbf{X}\in\mathbb{X}}\mathcal{F}(\mathbf{X})$ such that $\mathbf{X}\in\mathbb{D}$, where $\mathbb{D}$ is a convex set, can be solved using Lion-$\mathcal{K}$ with $\mathcal{K}^{*}(\mathbf{X})=\chi_{\mathbb{D}}(\mathbf{X})$. As in the cases of Lion and Muon, a particularly important instance is when $\mathbb{D}$ is a norm ball $\mathbb{B}_{\left\lVert\cdot\right\rVert}(\cdot)$. In this setting, $\nabla\mathcal{K}(\mathbf{X})$ corresponds to the solution of a linear minimization oracle problem $\max_{\mathbf{Z}\in\mathbb{X}}\left\langle\mathbf{X},\mathbf{Z}\right\rangle$ such that $\left\lVert\mathbf{Z}\right\rVert\leq 1$. Related discussions can be found, for example, in [BN24] and [PXA+25].
@@ 7 Convergence analysis of Muon
In this section, we generalize the analysis of [CLLL24] to handle $\mathbb{X}$-valued updates and leverage this result to provide convergence rates for the KKT score function (6) and prove the convergence of Muon to a KKT point of (5). Our strategy consists of three components:
@@
  • In Section 7.4, we put together the previous results and conclude that Muon converges to a KKT point of (5) using LaSalle’s invariance principle.
@@ 7.1 KKT points of spectral-norm-constrained problems
Then $\lambda\mathbf{X}^{\star}$ is a subgradient of $\left\lVert\cdot\right\rVert_{\textup{tr}}$ at $-\nabla\mathcal{F}(\mathbf{X}^{\star})$, since for all $\mathbf{Y}\in\mathbb{X}$,
@@
Let $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ be a singular value decomposition of $-\nabla\mathcal{F}(\mathbf{X}^{\star})$. By Fact 2,
@@
    where $r:=\operatorname{rank}(\nabla\mathcal{F}(\mathbf{X}^{\star}))$. In addition, $\left\lVert\mathbf{W}\right\rVert_{\textup{op}}\leq 1$ implies that $\boldsymbol{\sigma}_{1}(\mathbf{X}^{\star})=\dots=\boldsymbol{\sigma}_{r}(\mathbf{X}^{\star})=\frac{1}{\lambda}$, with corresponding singular vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{r}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{r}$. But by construction, $\nabla\mathcal{F}(\mathbf{X}^{\star})$ has singular values $\boldsymbol{\mu}$ and singular vectors $-\mathbf{u}_{1},\dots,-\mathbf{u}_{n}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{m}$. Thus
@@ 7.2 Convergence rate of Lion-$\mathcal{K}$ with deterministic gradient
Under Assumptions 1, 2, and 4, let $10<\beta_{1}<\beta_{2}<$ 1 and $\lambda>0$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3). Let
$c_{t}$ $:=\frac{\eta_{t}\lambda\beta_{1}}{\eta_{t}\lambda(1-\beta_{1})+(1-\beta_{2})}$ (12)
$b_{t}$ $:=\frac{\beta_{1}(1-\beta_{2})}{(\beta_{2}-\beta_{1})(\eta_{t}\lambda(1-\beta_{1})+(1-\beta_{2}))}$
$a_{t}$ $:=c_{t}+1$
$\mathcal{H}_{t}$ $:=\mathcal{F}(\mathbf{X}_{t})-\mathcal{F}^{\star}+\frac{1}{\lambda}\mathcal{K}^{*}(\lambda\mathbf{X}_{t})+\frac{c_{t}}{\lambda}(\mathcal{K}^{*}(\lambda\mathbf{X}_{t})+\mathcal{K}(\mathbf{M}_{t})-\left\langle\lambda\mathbf{X}_{t},\mathbf{M}_{t}\right\rangle)$
$\Gamma_{t}$ $:=\left\langle\nabla\mathcal{K}(\widetilde{\mathbf{M}}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\nabla\mathcal{K}^{*}(\lambda\mathbf{X}_{t+1})\right\rangle$
$\Delta_{t}$ $:=\left\langle\nabla\mathcal{K}(\widetilde{\mathbf{M}}_{t+1})-\nabla\mathcal{K}(\mathbf{M}_{t+1}),\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle.$
@@
$\phantom{=}+\frac{b_{t}\eta_{t}\lambda}{c_{t}}\left\langle\frac{c_{t}}{\lambda}\nabla\mathcal{K}(\mathbf{M}_{t+1})-c_{t}\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\phantom{=}+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\phantom{=}+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
[…]
@@
$\phantom{=}+\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\widetilde{\mathbf{M}}_{t}\right\rangle$
$\phantom{=}-\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rangle$
@@
$\phantom{=}-\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rangle$
@@
$\phantom{=}+\frac{2\beta_{1}C_{\mathcal{K}}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}}{(1-\beta_{2})T}+\frac{4\eta C_{\mathcal{K}}^{2}L(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}$
@@
Under Assumptions 1, 2, and 4, let $10<\beta_{1}<\beta_{2}<$ 1, $\lambda>0$, $\eta_{t}=\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$, and $\left\lVert\lambda\mathbf{X}_{0}\right\rVert_{\textup{op}}\leq 1$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3) with a deterministic gradient and that $\nabla\mathcal{K}$ has bounded norm. Then
@@
Moreover, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2) with a deterministic gradient,
@@
In the setting of Proposition 3, note that $\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$ and both $\nabla\mathcal{K}(\widetilde{\mathbf{M}})$ and $\lambda\mathbf{X}$ having bounded norm implies that the right-hand side of (13) is $O\left(\frac{1}{T}\right)$. The claim follows after dividing both sides by $\eta$ and realizing $\Gamma_{t},\Delta_{t}\geq 0$ by Lemma 1 and $a_{t},b_{t}\geq 0$.
@@ 7.3 Convergence rate of Lion-$\mathcal{K}$ with stochastic gradient
We now show results analogous to the ones in Section 7.2 when using (3) with a stochastic gradient. The following lemmas will be useful for bounding the noise arising from stochastic gradients.
@@
[…]
@@
[…]
@@
In the setting of Theorem 2 and under Assumption 3, let $\mathcal{K}$ be a norm on $\mathbb{X}$, and suppose instead that $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3) with a stochastic gradient. Then
@@
Moreover, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2) with a stochastic gradient,
@@
$\phantom{=}+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\phantom{=}+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\phantom{=}+2\beta_{1}C_{\mathcal{K}}\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\mathbb{E}[\left\lVert\mathbf{G}_{k}-\mathbf{G}_{k-1}\right\rVert_{\textup{F}}]+2\beta_{1}\beta_{2}^{t-1}C_{\mathcal{K}}\mathbb{E}[\left\lVert\mathbf{G}_{0}+\mathbf{M}_{0}\right\rVert_{\textup{F}}]$
@@
$\phantom{=}+\frac{2\beta_{1}C_{\mathcal{K}}}{1-\beta_{2}}\left(\frac{2\sigma}{\sqrt{n_{\mathrm{batch}}}}+2\eta C_{\mathcal{K}}L\right)+2\beta_{1}\beta_{2}^{t-1}C_{\mathcal{K}}\mathbb{E}[\left\lVert\mathbf{G}_{0}+\mathbf{M}_{0}\right\rVert_{\textup{F}}]$
@@
$\phantom{=}+\frac{4C_{\mathcal{K}}(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}\left(\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}+\eta C_{\mathcal{K}}L\right)$
@@
$\phantom{=}+\frac{4C_{\mathcal{K}}(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}\left(\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}+\eta C_{\mathcal{K}}L\right),$
@@
7.4 Convergence of Muon to a KKT point
In this section, we use LaSalle’s invariance principle, Proposition 1, and Theorem 3 to show that Muon converges to a KKT point of (5).

### v2

@@ 6.2 Generalizations of Muon
Quadratic form $\textup{Tr}(\mathbf{X}^{\top}\mathbf{M}\mathbf{X})$ All matrices, $\mathbf{M}\succeq\mathbf{0}$ Smooth quadratic form, convex
@@
More generally, the constrained optimization problem $\min_{\mathbf{X}\in\mathbb{X}}\mathcal{F}(\mathbf{X})$ such that $\mathbf{X}\in\mathbb{D}$, where $\mathbb{D}$ is a convex set, can be solved using Lion-$\mathcal{K}$ with $\mathcal{K}^{*}(\mathbf{X})=\chi_{\mathbb{D}}(\mathbf{X})$. As in the cases of Lion and Muon, a particularly important instance is when $\mathbb{D}$ is a norm ball $\mathbb{B}_{\left\lVert\cdot\right\rVert}(1)$. In this setting, $\nabla\mathcal{K}(\mathbf{X})$ corresponds to the solution of a linear minimization oracle problem $\max_{\mathbf{Z}\in\mathbb{X}}\left\langle\mathbf{X},\mathbf{Z}\right\rangle$ such that $\left\lVert\mathbf{Z}\right\rVert\leq 1$. Related discussions can be found, for example, in [BN24] and [PXA+25].
@@ 7 Convergence analysis of Muon
In this section, we generalize the analysis of [CLLL24] to handle $\mathbb{X}$-valued updates and leverage this result to provide convergence rates for the KKT score function (6) and prove the convergence of Muon to the set of KKT points of (5). Our strategy consists of three components:
@@
  • In Section 7.4, we put together the previous results and conclude that Muon converges to the set of KKT points of (5) using LaSalle’s invariance principle.
@@ 7.1 KKT points of spectral-norm-constrained problems
Then $\lambda\mathbf{X}^{\star}$ is a subgradient of $\left\lVert\cdot\right\rVert_{\textup{tr}}$ at $\nabla\mathcal{F}(\mathbf{X}^{\star})$, since for all $\mathbf{Y}\in\mathbb{X}$,
@@
Let $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ be a singular value decomposition of $\nabla\mathcal{F}(\mathbf{X}^{\star})$. By Fact 2,
@@
    where $r:=\operatorname{rank}(\nabla\mathcal{F}(\mathbf{X}^{\star}))$. In addition, $\left\lVert\mathbf{W}\right\rVert_{\textup{op}}\leq 1$ implies that $\boldsymbol{\sigma}_{1}(\mathbf{X}^{\star})=\dots=\boldsymbol{\sigma}_{r}(\mathbf{X}^{\star})=\frac{1}{\lambda}$, with corresponding singular vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{r}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{r}$. But by construction, $\nabla\mathcal{F}(\mathbf{X}^{\star})$ has singular values $\boldsymbol{\mu}$ and singular vectors $\mathbf{u}_{1},\dots,-\mathbf{u}_{n}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{m}$. Thus
@@ 7.2 Convergence rate of Lion-$\mathcal{K}$ with deterministic gradient
Under Assumptions 1, 2, and 5, let $0\leq\beta_{1}<\beta_{2}<1$ and $\lambda>0$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3). Let
$c_{t}$ $=\frac{\eta_{t}\lambda\beta_{1}}{\eta_{t}\lambda(1-\beta_{1})+(1-\beta_{2})}$ (12)
$b_{t}$ $=\frac{\beta_{1}(1-\beta_{2})}{(\beta_{2}-\beta_{1})(\eta_{t}\lambda(1-\beta_{1})+(1-\beta_{2}))}$
$a_{t}$ $=c_{t}+1$
$\mathcal{H}_{t}$ $=\mathcal{F}(\mathbf{X}_{t})-\mathcal{F}^{\star}+\frac{1}{\lambda}\mathcal{K}^{*}(\lambda\mathbf{X}_{t})+\frac{c_{t}}{\lambda}(\mathcal{K}^{*}(\lambda\mathbf{X}_{t})+\mathcal{K}(\mathbf{M}_{t})-\left\langle\lambda\mathbf{X}_{t},\mathbf{M}_{t}\right\rangle)$
$\Gamma_{t}$ $=\left\langle\nabla\mathcal{K}(\widetilde{\mathbf{M}}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\nabla\mathcal{K}^{*}(\lambda\mathbf{X}_{t+1})\right\rangle$
$\Delta_{t}$ $=\left\langle\nabla\mathcal{K}(\widetilde{\mathbf{M}}_{t+1})-\nabla\mathcal{K}(\mathbf{M}_{t+1}),\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle.$
@@
$\quad+\frac{b_{t}\eta_{t}\lambda}{c_{t}}\left\langle\frac{c_{t}}{\lambda}\nabla\mathcal{K}(\mathbf{M}_{t+1})-c_{t}\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\quad+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\quad+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
Let $\mathcal{K}(\mathbf{X})=\left\lVert\mathbf{X}\right\rVert$ for a norm $\left\lVert\cdot\right\rVert$ on $\mathbb{X}$. Then $\left\langle\nabla\mathcal{K}(\mathbf{X}),\mathbf{X}\right\rangle=\mathcal{K}(\mathbf{X})$.
@@
$\quad+\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\widetilde{\mathbf{M}}_{t}\right\rangle$
$\quad-\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rangle$
@@
$\quad-\left\langle-\operatorname{msgn}(\nabla\mathcal{F}(\mathbf{X}_{t}))-\operatorname{msgn}(\widetilde{\mathbf{M}}_{t}),\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rangle$
@@
$\quad+\frac{2\beta_{1}C_{\mathcal{K}}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}}{(1-\beta_{2})T}+\frac{4\eta C_{\mathcal{K}}^{2}L(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}$
@@
Under Assumptions 1, 2, and 5, let $0\leq\beta_{1}<\beta_{2}<1$, $\lambda>0$, $\eta_{t}=\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$, and $\left\lVert\lambda\mathbf{X}_{0}\right\rVert_{\textup{op}}\leq 1$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3) with deterministic gradients and that $\nabla\mathcal{K}$ has bounded norm. Then
@@
Moreover, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2) with deterministic gradients,
@@
[…]
@@ 7.3 Convergence rate of Lion-$\mathcal{K}$ with stochastic gradient
We now show results analogous to the ones in Section 7.2 when using (3) with stochastic gradients. The following lemmas will be useful for bounding the noise arising from stochastic gradients.
@@
Let $\mathcal{K}(\mathbf{X})=\left\lVert\mathbf{X}\right\rVert$ for a norm $\left\lVert\cdot\right\rVert$ on $\mathbb{X}$. Then for all $\mathbf{X}\in\mathbb{X}$, $\mathcal{K}^{*}(\nabla\mathcal{K}(\mathbf{X}))=0$.
@@
Let $\mathbf{X},\mathbf{Y}$ be $\mathbb{X}$-valued random variables satisfying $\textup{Var}(\mathbf{Y})<\infty$, and let $\mathcal{K}(\mathbf{X})=\left\lVert\mathbf{X}\right\rVert$ for a norm $\left\lVert\cdot\right\rVert$ on $\mathbb{X}$. Then there exists a constant $C_{\mathcal{K}}$ such that
@@
In the setting of Theorem 3 and under Assumption 3, let $\mathcal{K}$ be a norm on $\mathbb{X}$, and suppose instead that $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (3) with stochastic gradients. Then
@@
Moreover, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2) with stochastic gradients,
@@
$\quad+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\quad+b_{t}\eta_{t}\left\langle\nabla\mathcal{K}(\mathbf{M}_{t+1})-\lambda\mathbf{X}_{t+1},\widetilde{\mathbf{M}}_{t+1}-\mathbf{M}_{t+1}\right\rangle+\frac{L}{2}\left\lVert\mathbf{X}_{t+1}-\mathbf{X}_{t}\right\rVert_{\textup{F}}^{2}$
@@
$\quad+2\beta_{1}C_{\mathcal{K}}\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\mathbb{E}[\left\lVert\mathbf{G}_{k}-\mathbf{G}_{k-1}\right\rVert_{\textup{F}}]+2\beta_{1}\beta_{2}^{t-1}C_{\mathcal{K}}\mathbb{E}[\left\lVert\mathbf{G}_{0}+\mathbf{M}_{0}\right\rVert_{\textup{F}}]$
@@
$\quad+\frac{2\beta_{1}C_{\mathcal{K}}}{1-\beta_{2}}\left(\frac{2\sigma}{\sqrt{n_{\mathrm{batch}}}}+2\eta C_{\mathcal{K}}L\right)+2\beta_{1}\beta_{2}^{t-1}C_{\mathcal{K}}\mathbb{E}[\left\lVert\mathbf{G}_{0}+\mathbf{M}_{0}\right\rVert_{\textup{F}}]$
@@
$\quad+\frac{4C_{\mathcal{K}}(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}\left(\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}+\eta C_{\mathcal{K}}L\right)$
@@
$\quad+\frac{4C_{\mathcal{K}}(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}\left(\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}+\eta C_{\mathcal{K}}L\right),$
@@
7.4 Convergence of Muon to the set of KKT points
In this section, we use LaSalle’s invariance principle, Proposition 2, and Theorem 4 to show that Muon converges to the set of KKT points of (5).